ACCT 3000 – Assignment #1 – DUE Wednesday February 12, 2020

Kate and Claire, recent college graduates, are unable to find suitable jobs in their field of accounting.  However, each has been involved with a small business of their own for the last several years, and have been doing very well.   Kate is a talented seamstress, and has designed a line of fashionable blazers that are selling for $500/each.  Kate remains shocked at how fast the orders are coming in, and wonders if this could be something big.   Claire also has a small but growing business.  She manufactures synthetic leather belts that are selling for $50/each, and is similarly experiencing strong consumer interest.  A few large retailers have started to place orders with both girls, and both are  struggling to keep up with demand.  Kate and Claire are wondering if they should combine their lines  and start building the business together, since there is a high amount of overlap among their customers, and they could likely achieve some synergies by combining their marketing and customer service efforts. The belts go very well with the blazers. With a solid knowledge of their accounting basics, both have kept very thorough cost and marketing data.  So they decided to pull it all together and analyze it.

Kate’s blazers – manufacturing data

# of Blazers

Total Manufacturing Costs

2019

400

$140,000

2018

350

130,000

2017

310

122,000

2016

240

108,000

2015

275

115,000

2014

250

106,000

Kate’s blazers – marketing data

# of Blazers

Total Marketing Costs

2019

400

$60,000

2018

350

55,000

2017

310

51,000

2016

240

44,000

2015

275

47,500

2014

250

45,000

Claire’s belts – manufacturing data

# of Belts

Total Manufacturing Costs

2019

1,700

$66,500  

2018

1,400

56,000  

2017

1,100

45,500

2016

1,000

42,000

2015

1,200

49,000

2014

900

38,500

Claire’s belts – marketing data

# of Belts

Total Marketing Costs

2019

1,700

$11,500

2018

1,400

10,000

2017

1,100

8,500

2016

1,000

8,000

2015

1,200

9,000

2014

900

7,500

DELIVERABLE:

Prepare a comprehensive report with the answers to each of the following questions.  Provide all calculations in a well-organized manner, with the final answer for each part clearly stated in a complete sentence.  

Total report should be between 2-3 pages.  

QUESTIONS:

1.  High-low cost estimation method

  1. Use the high-low method to estimate the per-unit variable costs and total fixed costs for the blazers.
  2. Use the high-low method to estimate the per-unit variable costs and total fixed costs for the belts.

NOTE:  The variable and fixed costs for each product line, blazers and belts, contain both a manufacturing and a marketing component.  However, the data is recorded separately, which means FOUR separate high-low analyses must be conducted.

2. Cost-Volume-Profit (CVP) analysis, single-product

  1. Use CVP analysis to calculate the break-even point in units for the blazers.
  2. Use CVP analysis to calculate the break-even point in units for the belts.

3.  CVP, multiple-product setting

Merging the data together, it appears the sales mix is approximately 300 blazers and 1,200 belts each year.

For this CVP analysis, assume an additional $30,000 of combined fixed costs, this will be largely customer service costs.

  1. Calculate the break-even point for both product lines combined.

4.  Cost classification

  1. Classify the manufacturing costs, marketing costs, and customer service costs either as production expenses or period expenses.
  2. For the period expenses, further classify them into either selling expenses or general and administrative expenses.

5.  Sensitivity CVP analysis  and production versus period expenses – Multiple-Product Setting

  1. If both variable and fixed production expenses (refer to Question #1) associated with the blazer increased by 5% (beyond the estimate from the high-low analysis), how many blazers and belts would need to be sold in order to earn a target income of $96,000?  Assume the same sales mix and additional fixed costs in Question #3.

6.  Margin of Safety  

  1. Calculate the margin of safety (both in units and in sales dollars), assuming they sell 700 blazers and 2,500 belts next year.  Ignore the suggested change in expenses in Question #5.
  2. What does this margin of safety mean?

7.  Income statement

  1. Prepare a forecasted income statement (to the operating income line) for Kate and Claire, assuming they sell 700 blazers and 2,500 belts, and assuming the 5% cost increase for the blazer (variable and fixed production costs) and the additional $30,000 combined fixed costs.
  2. How would you suggest Kate and Claire divide the operating profit?

8.  Degree of Operating Leverage

  1. Calculate the degree of operating leverage, with the same assumptions as Question #7.
  2. What does the DOL tell Kate and Claire?  Do you think they should consider promotional events? Why or why not?
Answer #1

1) High low-cost estimation method

For Blazers

Variable cost per unit = Highest activity cost - Lowest activity cost / Highest activity units - Lowest activity units

   = (1,40,000+60,000) - (1,06,000+45,000) / 400 - 250

= 326.67 per unit

Fixed Cost = Highest activity cost - (Variable cost per units*Highest activity units)

= 1,40,000+60,000 - ( 326.67*400) = $ 69332

  OR

Fixed Cost = Lowest activity cost - ( Variable cost per units*Lowest activity units)

= 1,06,000+45,000 - (326.67*250)

= $ 69,332

For Belts

Variable cost per unit = Highest activity cost - Lowest activity cost / Highest activity units - Lowest activity units

= 66,500+11,500 - 38,500 + 7.500 / 1700 - 900

= $ 40 per unit

Fixed Cost = Highest activity cost - (Variable cost per units*Highest activity units)

= 66,500+11,500 - ( 40*1700) = $10,000

   OR

Fixed Cost = Lowest activity cost - ( Variable cost per units*Lowest activity units)

= 38500+7500 - ( 40*900) = $10000

2) Cost volume profit analysis

  For blazers

Selling price per unit = $500

Less: Variable cost per unit =   (326.67)

Contibution per unit $ 173.33

Break even point ( in units) = Fixed cost / contibution per unit = 69332 / 173.33 = 400   

For belts

Selling price per unit = $50

Less: Variable cost per unit =   (40)

Contribution per unit $ 10

Break even point ( in units) = Fixed cost / contibution per unit = $10000 / 10 = 1000

3) CPV multiple product setting

Combined contribution per unit = 173.33+10 = 183.33

Break even point for combined products line = Fixed cost / Contibution per unit

= 69332 +10000 +30000 / 183.33 = 596.37 units

4) a) Manufacutring cost = Production expenses

Marketing cost and customer service cost = Period cost

b) Marketing cost is a period cost.